For each list of integers, provide a simple formula or rule that generates the terms of an integer sequence that begins with the given list. Assume the sequence starts with . Do not forget initial conditions if required.

a)     7, 11, 15, 19, 23, 27, 31, 35, 39, 43

b)    -1, 2, 1, 3, 4, 7, 11, 18, 29, 47


1
Expert's answer
2021-11-22T14:36:20-0500

a)


"a_2-a_1=11-7=4,a_3-a_2=15-11=4,"

"a_4-a_3=19-15=4,a_5-a_4=23-19=4,"

"a_6-a_5=27-23=4,a_7-a_6=31-27=4,"

The sequence

"a_8-a_7=35-31=4,a_9-a_8=39-35=4,"

"a_{10}-a_9=43-39=4"

The sequence "\{a_n\}" is arithmetic progression with "a=7, d=4."


b)

"a_0=-1, a_1=2,"

"a_3=1=-1+2=a_1+a_2,"

"a_4=3=2+1=a_2+a_3,"

"a_5=4=1+3=a_3+a_4,"

"a_6=7=3+4=a_4+a_5,"

"a_7=11=4+7=a_5+a_6,"

"a_8=18=7+11=a_6+a_7,"

"a_9=29=11+18=a_7+a_8,"

"a_{10}=47=18+29=a_8+a_9"

The sequence "\{a_n\}" satisfies the recurrence relation "a_n=a_{n-2}+a_{n-1}, n=2, 3, ..." with "a_0=-1, a_1=2."